Find the gravitational force between the sun with a mass of 2 x 1030 kg and the Earth with a mass of 6 

x 1024 kg if the distance between them is 1.4 x 108 km


1
Expert's answer
2021-10-25T09:56:40-0400
"F_{grav}=\dfrac{GM_{S}M_E}{r_{SE}^2},""F_{grav}=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times2\times10^{30}\ kg\times6\times10^{24}\ kg}{(1.4\times10^{11}\ m)^2},""F_{grav}=4.1\times10^{22}\ N."